number theory and discrete math

Find Gaussian integers a+ib and r+is such that 234212 + 3421 i - (23+ 41i) (a + bi) + (r + si) such that 2 (72 +82) < 232 + 4show full steps
Answer #1

234212 + 3421() = (23+41 (atbi) + (2+1) 23a+23bit4101-416 totsi - (238 – 416+r) +i (23b+4lats 2341 -41b+7= 234212 and 23b +4lfrom eg.) 2210 + 23(?) tulcea =5527139. 2210a- 5525574.294 ao 2500 from ego, 2.3(2500) - 4164 23 =234212 r 416-723(2500)+22 -

I hope explaination is clear.

Comment for doubts and upvote if u like it.

Thank you.